# Fixing Code Generation Errors for Large Language Models

### v1

@@ III-C1 Motivation and Method
    Phind, was introduced by the Phind team, available in two versions [52]: Phind-CodeLlama-34B-v1 and Phind-CodeLlama-34B-v2. According to the team’s report [45], the latter version even outperforms OpenAI’s GPT series in code generation.
  • $\bullet$
    WizardCoder, was introduced by Microsoft [40], which empowers code generation LLMs with complex instruction fine-tuning. We re-implemented the versions of WizardCoder-Python in different parameter sizes (7B/13B/34B) [53].
  • $\bullet$
    CodeLlama, was introduced by Meta AI [19], which is a competitive code generation model and inspired many other varied versions such as CodeTulu [47] and WizardCoder [40]. We used the CodeLlama-python (7B/13B/34B) [54] as our baseline, which fine-tunes CodeLlama with more Python code and shows better performance in code generation.
  • $\bullet$
    Phi, was released by Microsoft [46], which is renowned for its smaller model size (1.3B) but demonstrated exceptional high performance in code generation. Later, Microsoft released new versions including Phi-1.5 Li et al. [55] and Phi-2 [56]. Thus, these three models were selected as our baselines including Phi-1 (1.3B), Phi-1.5 (1.3B), and Phi-2 (2.7B) [57].
  • $\bullet$
    CodeTulu, was proposed by AllenAI [58], whose coding capabilities matched or even exceeded the performance of GPT-3.5-turbo in several benchmark tests. We selected the best-performing and variably parameterized versions of CodeTulu-2 (7B/13B/34B) as our baselines.
III-C2 LLMs Reproduction and Evaluation
Fig. 2 shows the accuracy distribution of 14 LLMs over ten times tests, sorted by the average accuracy from lowest to highest. The best-performing model is the Phind-CodeLlama-V2-34B, with an average accuracy of 63.29%. Table I lists the specific numerical characteristics of these 14 LLMs from the ten tests, including minimum, maximum, mean, standard deviation, and results reported in related research papers [9]. From the study by Zheng et al. [9], we know that the 14 LLM accuracies range from 34.10% to 73.80%, while our replicated test results have an average accuracy ranging from 27.14% to 63.29%. Compared to reported data, our average test results are approximately reaching 84.07% of the reported results effectively. This difference is within an acceptable range, likely related to the numerical precision used in the experiments. Additionally, the standard deviation (STD) for all models is significantly lower than the mean (less than 10% of the mean), indicating low data variability.
Figure 2: The distribution of test results for 14 models, sorted by the mean values from lowest to highest, where the thin vertical lines represent the means; the thick short lines inside the boxes represent the medians; and the stars indicate the reported performance in the related papers [9].
TABLE I: Accuracy of ten tests of LLMs on HumanEval dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values, compared to the result reported in related papers [9], where the percentage value in parenthesis indicates the improvement ratio to the reported; ”-” indicates the related paper did not report the Pass@1 as required in our study but Pass@10.
@@ III-C2 LLMs Reproduction and Evaluation
CodeLlama-Python-7B 38.40 31.10 (--19.01%) 35.37 (--07.89%) 32.69 (--14.87%) 1.39
CodeLlama-Python-13B 43.30 31.71 (--26.77%) 40.24 (--07.07%) 36.65 (--15.36%) 2.82
CodeLlama-Python-34B 53.70 39.02 (--27.34%) 46.34 (--13.71%) 43.72 (--18.58%) 2.17
WizardCoder-Python-V1.0-7B 55.50 40.05 (--27.75%) 43.90 (--20.90%) 42.23 (--23.91%) 1.05
WizardCoder-Python-V1.0-13B 64.00 48.17 (--24.73%) 56.10 (--12.34%) 52.13 (--18.54%) 2.48
WizardCoder-Python-V1.0-34B 73.20 57.32 (--21.72%) 62.80 (--14.21%) 59.09 (--19.29%) 1.58
Phind-CodeLlama-V2-34B 73.80 60.37 (--18.20%) 65.24 (--11.60%) 63.29 (--14.24%) 1.51
Phi-1-1.3B 50.60 43.29 (--14.45%) 46.95 (--07.21%) 45.30 (--10.47%) 1.04
Phi-1.5-1.3B 34.10 23.78 (--30.26%) 28.66 (--15.95%) 27.14 (--20.41%) 1.42
@@
To assess reproducibility, we calculated the Pearson correlation coefficient [59] between our mean values and the reported values. The correlation coefficient of about $0.989$ indicates almost perfect consistency between our results and the reported data. The corresponding P-value of $6.91\times 10^{-9}$ is highly significant statistically, demonstrating a very high level of replication. Phi-2 and the CodeTulu-2 family of LLMs did not report the usual Pass@1 values. Instead, the CodeTulu-2 family chose to report Pass@10 to highlight the superiority of their models. In our study, these models had an average accuracy of 39.26%, ranking ninth among all models.
Table II shows the numerical characteristics of inference times per sample for 13 types of open-source LLMs mentioned, over ten tests, including minimum, maximum, average, and standard deviation values. It is worth noting that GPT-3.5-turbo, being a closed-source model running on OpenAI’s servers, operates in a different testing environment from the other 13 open-source models. Consequently, we have excluded its inference time from this analysis. From the test results, the average inference times ranged from 16.84 seconds to 77.38 seconds, with the LLM Phi-1-1.3B having the shortest inference time and WizardCoder-Python-V1.0-34B having the longest. Additionally, the standard deviations (STD) of average inference times for most LLMs were significantly lower than their corresponding averages (less than 10% of the mean), indicating low variability in inference times across most models.
TABLE II: The average testing time used per sample in ten LLM tests conducted on the HumanEval dataset with a total of 164 samples, in terms of the minimum, maximum, mean, and standard deviation (STD).
Model Min Max Mean STD
CodeLlama-Python-7B 25.18s 29.26s 27.33s 1.10
CodeLlama-Python-13B 31.40s 48.73s 41.42s 4.46
CodeLlama-Python-34B 65.29s 82.66s 76.49s 6.02
WizardCoder-Python-V1.0-7B 21.98s 25.45s 24.03s 1.12
WizardCoder-Python-V1.0-13B 17.16s 19.93s 18.12s 0.87
WizardCoder-Python-V1.0-34B 72.94s 82.53s 77.38s 2.66
Phind-CodeLlama-V2-34B 24.73s 29.43s 26.91s 1.16
Phi-1-1.3B 16.33s 17.28s 16.84s 0.25
Phi-1.5-1.3B 16.27s 17.97s 17.14s 0.55
Phi-2-2.7B 21.39s 22.72s 22.18s 0.36
CodeTulu-2-7B 33.27s 34.49s 33.96s 0.35
CodeTulu-2-13B 38.40s 56.50s 47.23s 6.23
CodeTulu-2-34B 22.90s 30.10s 25.37s 2.18
III-C3 Impact Analysis
Fig. 3 presents a linear regression analysis of the impact of parameter size on both model performance and average inference times. Through this analysis, the figure demonstrates two key relationships. On the model performance, the regression equation $y=0.5772x+34.083$ with a goodness of fit $R^{2}=0.5379$ reveals that, generally, the larger the model’s scale, the better its performance. The moderate goodness of fit value suggests that this statistical result has a certain level of significance. The data analysis from Table I similarly shows that this trend applies to most LLMs. To further analyze the performance of LLMs with different parameter sizes, we presented a bar graph, Fig. 4, featuring four groups of LLMs. Among LLMs with more than 20B parameters, the top three performers—Phind-CodeLlama-V2-34B, WizardCoder-Python-V1.0-34B, and GPT-3.5-turbo—demonstrated comparable performance. For LLMs with parameters between 10B and 20B, the best performer was WizardCoder-Python-V1.0-13B. In the range of 5B to 10B parameters, the best performer was WizardCoder-Python-V1.0-7B. For LLMs with less than 5B parameters, the best was Phi-1-1.3B. We observed that some LLMs, despite their smaller scale, achieved or even surpassed the accuracy of larger-scale LLMs. For instance, Phi-1-1.3B, though small in scale, achieved an accuracy of 45.3%, exceeding that of the 34B Codellama-Python model, which had an accuracy of 43.72%. This demonstrates that the size of the model’s parameters is not the sole determinant of performance.
Figure 3: Scatter plot and linear fit curve of accuracy versus inference time for 13 LLMs.
(a) Accuracy of models with less than 5B parameters.
(b) Accuracy of models with parameters between 5B and 10B.
(c) Accuracy of models with parameters between 10B and 20B.
(d) Accuracy of models with more than 20B parameters.
Figure 4: Bar chart of accuracy for four groups of models
On the average inference times, the regression function $y=0.9496x+20.249$ with a goodness of fit $R^{2}=0.3806$ is displayed. Compared to the fitting for accuracy, this $R^{2}$ value is significantly lower, suggesting substantial random fluctuations in the data and weak explanatory power of the regression function for inference times. This is confirmed by the data in Table II. For example, Phind-CodeLlama-V2-34B and CodeTulu-2-34B average about 25 to 26 seconds, while the smaller-scale CodeLlama-Python-13B requires 41.42 seconds. This is speculated to be related to the length of content generated by the models. Observing the file sizes of model outputs, we found that the former models’ output files are approximately 2MB, compared to 3.3MB for the latter. This indicates that the former models can solve the same problem with shorter code. Consequently, the time required to generate this shorter content is also shorter, resulting in a shorter overall inference time.
@@ III-C3 Impact Analysis
[nobreak=true] Answer to RQ1: All our reproduced LLMs exhibit close performance as the original reports. The accuracy of LLMs is positively correlated with the size of their parameters in general except for two small-sized LLMs: Phi-1-1.3B and WizardCoder-Python-V1.0-7B. The average inference time of LLMs is not correlated with LLM scale, large LLMs such as Phind-CodeLlama-V2-34B and CodeTulu-2-34B demonstrate significantly shorter inference times than smaller-scale LLMs.
III-D Why did These LLMs Fail in Code Generation? (RQ2)
III-D1 Motivation and Method
To investigate why these LLMs failed in code generation, as described in RQ2, we extracted the error information from failed test cases and developed a systematic process for error analysis. The HumanEval dataset consists of 164 code generation tasks, each accompanied by an average of 7.7 corresponding test cases [13]. Our objective is to take the generated code and test cases as input and return the test results, assessing whether the code passed or failed. Utilizing the model and dataset interfaces provided by HuggingFace, we obtained the code generated by the model for each sample in the dataset, along with the corresponding test cases, all presented in string format. These strings can be parsed and executed as a series of Python statements using Python’s built-in exec() function, thus enabling automated testing. If the generated code passes all test cases, the test continues running; otherwise, the generated code fails one or more test cases, reports an exception, and terminates.
@@ III-D1 Motivation and Method
Figure 5: After testing each sample in the dataset, the output file displays the structure of the samples, including an example of a sample that passed the test and another that did not pass the test.
III-D2 Error Classification and Causes Analysis
Through ten rounds of testing on 14 different LLMs using the HumanEval dataset, we collected a total of 12,837 errors. Based on the Test Result Information shown in Fig. 5, we categorized these errors into 14 Python exception types. Fig. 6 presents a three-dimensional bar chart illustrating the frequency of occurrence for each of these 14 error types across all LLMs over the ten test rounds. Moreover, considering that the same type of error can be triggered by various causes, we unanimously agreed that it is necessary not only to categorize errors but also to analyze the code containing errors, thereby identifying one or more causes leading to each type of error. The above research results are summarized in Table III. The following is a detailed explanation of each error type:
Figure 6: Distribution of total errors across 14 LLMs in ten test runs.
TABLE III: Error types and causes summary: categories, frequencies, causes, and code examples. Note that examples are provided in Appendix.
@@ III-D2 Error Classification and Causes Analysis
III-D3 Distribution of Error Types
@@
Based on Fig. 4 and Table III, We have also found that the 14 types of errors can be categorized into three groups based on their frequency of occurrence, as illustrated in Fig. 7:
@@
Figure 7: Taxonomy of errors introduced while generating code using LLM.
From the perspective of errors, within the category of “The most frequent errors”, AssertionError and NameError exhibit distinctly different trends: NameError is relatively evenly distributed while the frequency of AssertionError significantly decreases as the model’s accuracy improves. To verify this, we calculated the Pearson correlation coefficients between the number of these two types of errors and the corresponding model’s accuracy. The results reveal a strong negative correlation between AssertionError occurrences and model accuracy, with a Pearson correlation coefficient of -0.84 and a highly significant p-value of $1.7\times 10^{-4}$. This suggests that as model accuracy improves, the frequency of AssertionErrors decreases substantially. In contrast, NameError shows a weaker negative correlation with model accuracy, having a correlation coefficient of -0.41 and a non-significant p-value of 0.14. This indicates that while there is a slight tendency for NameErrors to decrease as accuracy improves, the relationship is not statistically significant and the trend is less pronounced compared to AssertionErrors. For ”Occasionally occurring errors,” we also conducted a similar correlation analysis and found that SyntaxError and IndexError do not show a significant correlation, while the Pearson correlation coefficients for TypeError and ValueError are around -0.7, indicating a strong negative correlation. Since errors in the “Rarely occurring errors” category hardly ever occur, we could not conduct more in-depth statistical analysis on them.
From the model’s perspective, we found that only GPT-3.5-turbo and Phi-2-2.7B encountered ModuleNotFoundError, indicating that they not only called modules from the Python standard library but might also have used modules from non-standard libraries to aid in problem-solving. Additionally, GPT-3.5-turbo had the fewest number of AssertionErrors and the highest number of NameErrors, mainly because it often calls modules directly without adding import statements. This point is a significant consideration in our proposed solution, indicating that our approach is likely to be very effective in this model. Moreover, we also found that GPT-3.5-turbo and Phind-CodeLlama-V2-34B rarely experience SyntaxError, possibly because they are more proficient in mastering Python syntax and generally do not generate unnecessary functions, thus avoiding SyntaxErrors caused by incomplete functions. In contrast, Phi-1.5-1.3B and Phi-2-2.7B exhibited a higher frequency of SyntaxError occurrences. This could be attributed to their tendency to generate more complex or convoluted code structures, leading to a greater likelihood of syntax-related issues such as incomplete or incorrect function definitions.
III-D4 The relationship between failed samples and errors
@@
A single piece of code may contain multiple errors. However, the Analysis Process described in III-D is currently limited to extracting only the first error information from the code. To address this limitation, we propose a manual analysis process, illustrated in Fig. 8. This process aims to identify all errors in the code through an iterative “fix-test” cycle. By applying this process to code generated by LLMs, we confirmed that a single code snippet can contain multiple errors, establishing a one-to-many relationship between failed code snippets and errors. For example, the prompt from the HumanEval dataset (task_id 133), which serves as the comment in Listing LABEL:list_ft1, specifies the task for the LLM: squaring each number in a given list of digits and rounding each element to the upper integer (ceiling) first. The original code generated by the model, shown in Listing LABEL:list_ft1, triggered a SyntaxError due to an unclosed “[”. Based on the error type and the information provided, we manually corrected the error, resulting in the code shown in Listing LABEL:list_ft2. Subsequent testing revealed that this corrected code still failed to pass all tests. The error occurred because a generator expression was passed into the int() function, which expects a string, bytes, or a real number, leading to a TypeError. Consequently, we performed a second manual error correction, resulting in the code shown in Listing LABEL:list_ft3. After this final correction, the code successfully passed all test cases.
Figure 8: Schematic Diagram of the “Fix-Test” Process.
@@
[nobreak=true] Answer to RQ2: We collected a total of 12,837 errors, which can be classified into 14 categories according to Python’s built-in exception types. Among these, AssertionError emerged as the most common error type. The causes of these errors can be attributed to 19222The “Error Cause” column in Table III has 20 rows, but both ValueError and NotImplementError can be caused by the same cause, “Intentional Raise”. Therefore, there are only 19 distinct error causes in total. different causes, such as incorrect function implementation, syntax mistakes, logical errors, and failure to handle edge cases.
III-E Which Errors Can be Directly Fixed? (RQ3)
III-E1 Motivation and Method
@@ 3. $\bullet$
TABLE IV: Analysis of error causes for automatic fixability based on criteria mentioned in subsection III-E
@@
      Empty Function       ✓       ✓       $\times$       $\times$
@@
      Missing Content       ✓       ✓       $\times$       $\times$
      Missing Import       ✓       ✓       ✓       ✓
      Unbalanced Delimiters       $\times$       ✓       $\times$       $\times$
      Function Overflow       ✓       ✓       ✓       ✓
      Empty Sequence       $\times$       ✓       $\times$       $\times$
      Intentional Raise       $\times$       ✓       $\times$       $\times$
      Inappropriate Argument       $\times$       ✓       $\times$       $\times$
      Out of Bounds       $\times$       ✓       $\times$       $\times$
@@
      Non-existent Attribute       $\times$       ✓       $\times$       $\times$
@@
      Inconsistent Indentation       ✓       ✓       ✓       ✓
      Missing Module       $\times$       ✓       ✓       $\times$
@@
III-E2 Error Analysis Results
Based on the above criteria, we analyzed the fixability of each error type and its corresponding reasons. The results are summarized in Table IV. A detailed explanation of this table is as follows:

### v2

@@ 3.3.1. Motivation and Method
    CodeTulu, was proposed by AllenAI (AllenAI, 2023), whose coding capabilities matched or even exceeded the performance of GPT-3.5-turbo in several benchmark tests. We selected the best-performing and variably parameterized versions of CodeTulu-2 (7B/13B/34B) as our baselines.
3.3.2. LLMs Reproduction and Evaluation
Fig. 1 shows the accuracy distribution of 14 LLMs over ten times tests, sorted by the average accuracy from lowest to highest. The best-performing model is the Phind-CodeLlama-V2-34B, with an average accuracy of 63.29%. Table 1 lists the specific numerical characteristics of these 14 LLMs from the ten tests, including minimum, maximum, mean, standard deviation, and results reported in related research papers (Zheng et al., 2023). From the study by Zheng et al. (Zheng et al., 2023), we know that the 14 LLM accuracies range from 34.10% to 73.80%, while our replicated test results have an average accuracy ranging from 27.14% to 63.29%. Compared to reported data, our average test results are approximately reaching 84.07% of the reported results effectively. This difference is within an acceptable range, likely related to the numerical precision used in the experiments. Additionally, the standard deviation (STD) for all models is significantly lower than the mean (less than 10% of the mean), indicating low data variability.
Figure 1. The distribution of test results for 14 models, sorted by the mean values from lowest to highest, where the thin vertical lines represent the means; the thick short lines inside the boxes represent the medians; and the stars indicate the reported performance in the related papers (Zheng et al., 2023).
Table 1. Accuracy of ten tests of LLMs on HumanEval dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values, compared to the result reported in related papers (Zheng et al., 2023), where the percentage value in parenthesis indicates the improvement ratio to the reported; “-” indicates the related paper did not report the Pass@1 as required in our study but Pass@10.
@@ 3.3.2. LLMs Reproduction and Evaluation
CodeLlama-Python-7B 38.40 31.10 ($\downarrow$19.01%) 35.37 ($\downarrow$07.89%) 32.69 ($\downarrow$14.87%) 1.39
CodeLlama-Python-13B 43.30 31.71 ($\downarrow$26.77%) 40.24 ($\downarrow$07.07%) 36.65 ($\downarrow$15.36%) 2.82
CodeLlama-Python-34B 53.70 39.02 ($\downarrow$27.34%) 46.34 ($\downarrow$13.71%) 43.72 ($\downarrow$18.58%) 2.17
WizardCoder-Python-V1.0-7B 55.50 40.05 ($\downarrow$27.75%) 43.90 ($\downarrow$20.90%) 42.23 ($\downarrow$23.91%) 1.05
WizardCoder-Python-V1.0-13B 64.00 48.17 ($\downarrow$24.73%) 56.10 ($\downarrow$12.34%) 52.13 ($\downarrow$18.54%) 2.48
WizardCoder-Python-V1.0-34B 73.20 57.32 ($\downarrow$21.72%) 62.80 ($\downarrow$14.21%) 59.09 ($\downarrow$19.29%) 1.58
Phind-CodeLlama-V2-34B 73.80 60.37 ($\downarrow$18.20%) 65.24 ($\downarrow$11.60%) 63.29 ($\downarrow$14.24%) 1.51
Phi-1-1.3B 50.60 43.29 ($\downarrow$14.45%) 46.95 ($\downarrow$07.21%) 45.30 ($\downarrow$10.47%) 1.04
Phi-1.5-1.3B 34.10 23.78 ($\downarrow$30.26%) 28.66 ($\downarrow$15.95%) 27.14 ($\downarrow$20.41%) 1.42
@@
To assess reproducibility, we calculated the Pearson correlation coefficient (Pearson, 1895) between our mean values and the reported values. The correlation coefficient of about $0.989$ indicates almost perfect consistency between our results and the reported data. The corresponding P-value of $6.91\times 10^{-9}$ is highly significant statistically, demonstrating a very high level of replication. Phi-2 and the CodeTulu-2 family of LLMs did not report the usual Pass@1 values. Instead, the CodeTulu-2 family chose to report Pass@10 to highlight the superiority of their models. In our study, these models had an average accuracy of 39.26%, ranking ninth among all models.
@@
[nobreak=true] Answer to RQ1: All our reproduced foundation LLMs exhibit close performance as the original reports: compared to reported data, our average test results are reaching 84.1% of the reported results effectively.
3.4. Why did These LLMs Fail in Code Generation? (RQ2)
3.4.1. Motivation and Method
To investigate why these LLMs failed in code generation, we extracted the error information from failed test cases and developed a systematic process for error analysis. The HumanEval dataset consists of 164 code generation tasks, each accompanied by an average of 7.7 corresponding test cases (Chen et al., 2021). Our objective is to take the generated code and test cases as input and return the test results, assessing whether the code passed or failed. Utilizing the model and dataset interfaces provided by HuggingFace, we obtained the code generated by the model for each sample in the dataset, along with the corresponding test cases, all presented in string format. These strings can be parsed and executed as a series of Python statements using Python’s built-in exec() function, thus enabling automated testing. If the generated code passes all test cases, the test continues running; otherwise, the generated code fails one or more test cases, reports an exception, and terminates.
@@ 3.4.1. Motivation and Method
Figure 2. After testing each sample in the dataset, the output file displays the structure of the samples, including an example of a sample that passed the test and another that did not pass the test.
3.4.2. Error Classification and Causes Analysis
Through ten rounds of testing on 14 different LLMs using the HumanEval dataset, we collected a total of 12,837 errors. Based on the Test Result Information shown in Fig. 2 which are defined and reported by compiler, we collected these errors for a total of 14 Python exception types. Fig. 3 presents a three-dimensional bar chart illustrating the frequency of occurrence for each of these 14 error types across all LLMs over the ten test rounds. Moreover, considering that the same type of error can be triggered by various causes, we unanimously agreed that it is necessary not only to categorize errors but also to analyze the code containing errors, thereby identifying one or more causes leading to each type of error. The above research results are summarized in Table 2. The following is a detailed explanation of each error type:
Figure 3. Distribution of total errors across 14 LLMs in ten test runs.
Table 2. Error types and causes summary: categories, frequencies, causes, and code examples. Note that examples are provided in the Appendix.
@@ 3.4.2. Error Classification and Causes Analysis
3.4.3. Distribution of Error Types
@@
Based on Table 2, we have found that the 14 types of errors can be categorized into three groups based on their frequency of occurrence. The groups and descriptions of the 14 types of errors are illustrated in Fig. 4.
@@
Figure 4. Taxonomy of errors introduced while generating code using LLM.
3.4.4. The relationship between failed samples and errors
@@
A single piece of code may contain multiple errors. However, the Analysis Process described in 3.4 is currently limited to extracting only the first error information from the code. To address this limitation, we apply a “fix-test” cycle that continually fixes the erroneous code until it passes all the test cases. By applying this cycle to code generated by LLMs, we confirmed that a single code snippet can contain multiple errors, establishing a one-to-many relationship between failed code snippets and errors. For example, the prompt from the HumanEval dataset (task_id 133), which serves as the comment in Listing LABEL:list_ft1, specifies the task for the LLM: squaring each number in a given list of digits and rounding each element to the upper integer (ceiling) first. The original code generated by the model, shown in Listing LABEL:list_ft1, triggered a SyntaxError due to an unclosed “[”. Based on the error type and the information provided, we manually corrected the error, resulting in the code shown in Listing LABEL:list_ft2. Subsequent testing revealed that this corrected code still failed to pass all tests. The error occurred because a generator expression was passed into the int() function, which expects a string, bytes, or a real number, leading to a TypeError. Consequently, we performed a second manual error correction, resulting in the code shown in Listing LABEL:list_ft3. After this final correction, the code successfully passed all test cases.
@@
[nobreak=true] Answer to RQ2: We collected a total of 12,837 errors, which can be classified into 14 categories according to Python’s built-in exception types. Among these, AssertionError emerged as the most common error type. The causes of these errors can be attributed to 19222The “Error Cause” column in Table 2 has 20 rows, but both ValueError and NotImplementError can be caused by the same cause, “Intentional Raise”. Therefore, we have 19 distinct error causes in total. different causes, such as incorrect function implementation, syntax mistakes, logical errors, and failure to handle edge cases.
3.5. Which Errors Can be Directly Fixed? (RQ3)
3.5.1. Motivation and Method
@@
Table 3. Analysis of error causes for automatic fixability based on criteria mentioned in subsection 3.5
@@
Empty Function $\times$ $\times$
@@
Missing Content $\times$ $\times$
Missing Import
Unbalanced Delimiters $\times$ $\times$ $\times$
Function Overflow
Empty Sequence $\times$ $\times$ $\times$
Intentional Raise $\times$ $\times$ $\times$
Inappropriate Argument $\times$ $\times$ $\times$
Out of Bounds $\times$ $\times$ $\times$
@@
Non-existent Attribute $\times$ $\times$ $\times$
@@
Inconsistent Indentation
Missing Module $\times$ $\times$
@@
3.5.2. Error Analysis Results
Based on the above criteria, we analyzed the fixability of each error type and its corresponding reasons. The results are summarized in Table 3. A detailed explanation of this table is as follows: